OpenStudy (anonymous):

Find approximate solutions for the equation in the interval [-pi,pi]: tan x = 3 csc x

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

i got to \(cos^2x+3cos~x-1=0\) im stuck after that

hero (hero):

\[\frac{\sin x}{\cos x} = \frac{3}{\sin x}, [-\pi, \pi]\]

OpenStudy (anonymous):

yes

hero (hero):

sin^2x - 3cos x = 0 (1 - cos^2x) - 3cos x = 0

OpenStudy (anonymous):

yes and that can be turned into cos^2(x)+3cos(x)-1=0

hero (hero):

Okay, so all you have to do is let cos(x) = y

hero (hero):

y^2 + 3y - 1 = 0

hero (hero):

Then factor

OpenStudy (anonymous):

dont you factor?

OpenStudy (anonymous):

there we go, that's what i was missing

hero (hero):

Your study guide doesn't tell you what to do?

OpenStudy (anonymous):

nope, just questions

hero (hero):

Well, that's pointless. Why would they not show you?

OpenStudy (anonymous):

no the teacher gave a semester exam study guide with only questions for us to practice ...she explained briefly but i forgot what she had done cuz i wasnt writing :P

OpenStudy (anonymous):

k so i did quadratic and got\[x={{3\pm\sqrt{5}}\over2}\]

OpenStudy (anonymous):

i meant \(x=\large{{{3\pm\sqrt{13}}\over2}}\)

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@rajathsbhat

OpenStudy (anonymous):

@hartnn

hero (hero):

You need to replace y with cos(x)

OpenStudy (anonymous):

yea i did....and then i did quadratic equation

hero (hero):

No, I mean replace it back

OpenStudy (anonymous):

\[cos~x={{3\pm\sqrt{5}}\over2}?\]

hero (hero):

\[\cos(x) = {{3\pm\sqrt{5}}\over2}\]

OpenStudy (anonymous):

\[cos~x={{3\pm\sqrt{13}}\over2}\]

OpenStudy (anonymous):

okay then what?

hero (hero):

Isolate x

OpenStudy (anonymous):

\[x=cos^{-1}\left({{3\pm\sqrt{13}}\over2}\right)?\]

hero (hero):

Yes, but one result will work, the other won't

OpenStudy (anonymous):

im supposed to get \(\pm1.263\) and i got 1.878 :S

OpenStudy (anonymous):

@hartnn input?

OpenStudy (anonymous):

i got